OpenStudy (baconfuntimes):

Will medal! Hannah is making chocolate chip cookies. The amount of brown sugar needed for each batch is 1 1/2. How many batches of cookies can she make if she has 7 1/2 cups of brown sugar?

OpenStudy (anonymous):

5 maybe, I did 7 1/2 divided by 1 1/2 and got 5, so yeah, 5.

OpenStudy (alexandervonhumboldt2):

first you need to convert these fraction to improper fractions and then divide. \[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\] \[1 \frac{ 1 }{ 2 }=\frac{ 3 }{ 2 }\] \[7\frac{ 1 }{ 2 }=\frac{ 15 }{ 2 }\] now divide:

OpenStudy (alexandervonhumboldt2):

\[\frac{ 15 }{ 2 }/\frac{ 3 }{ 2 }=\frac{ 15 }{ \cancel 2 }*\frac{ \cancel 2 }{ 3 }=5\]